# Shot noise in Aharonov-Bohm interferometers: Comparison of helical and conventional setups

### v1

@@ III.3.3 Limiting cases
$\mathcal{F}^{\text{HI}}=\tfrac{1}{2}\left[{1-({\mathcal{T}}^{\text{HI}})^{2}}\right].$ (42)
@@
The formulas for the conductance and Fano factor in case of different tunneling contacts can also be obtained in closed form and are presented in Appendix C.2. An analog of Eq. (42) in case of different tunneling contacts, in contrast to CI case, cf. Eq. (29) looks a bit more cumbersome
$\mathcal{F}^{\rm HI}=\frac{\left[1-({\mathcal{T}}^{\rm HI})^{2}\right]\sqrt{\sinh^{2}(2\beta)+({\mathcal{T}}^{\rm HI})^{2}}}{\sqrt{\sinh^{2}(2\beta)+({\mathcal{T}}^{\rm HI})^{2}}+{\mathcal{T}}^{\rm HI}\cosh(2\beta)}\,,$ (43)
with the asymmetry parameter $\beta=\tfrac{1}{2}\ln(t_{\text{R}}/t_{\text{L}})$, Eq. (36).
Figure 10: The combined dependence of Fano factor, $\mathcal{F}^{\text{HI}}$, and the transmission coefficient, $\mathcal{T}^{\text{HI}}$, on the transparency of the contact, $\lambda$, for different strength of the magnetic impurity, $R_{\theta}$, and the magnetic flux $\phi=0.4$, according to Eqs.  (37), (38). The black dashed line borders the area of maximum transmission coefficient for given $R_{\theta}$.
@@
$\mathcal{T}^{\rm HI}=\tfrac{1}{2}{\tanh[2\lambda]},\quad\mathcal{F}^{\rm HI}=\frac{1}{2\cosh^{2}[2\lambda]},$ (44)
$\mathcal{F}^{\rm HI}=\tfrac{1}{2}-2\left(\mathcal{T}^{\rm HI}\right)^{2}.$
@@
$\mathcal{T}^{\text{HI}}=1-\frac{R_{\theta}}{2},$ (45)
$\mathcal{F}^{\text{HI}}=R_{\theta}\frac{1-R_{\theta}}{2-R_{\theta}}=\frac{(2\mathcal{T}^{\text{HI}}-1)(1-\mathcal{T}^{\text{HI}})}{\mathcal{T}^{\text{HI}}}\,.$
@@
Equation (45) has a simple physical meaning. For $r=1,$ the windings are absent, and there are two parallel conductors, i.e. lower and upper arms of the HI. The lower arm, with spin $\downarrow$, is an ideal conductor with the transmission coefficient $\mathcal{T}_{\downarrow}=1$ while the transmission through the upper arm is controlled by the magnetic defect, $\mathcal{T}_{\uparrow}=\cos^{2}\theta,$ varying from 00 to $1$ Then, we get
@@
Taking into account that the total transmission coefficient in this case is given by $\mathcal{T}^{\text{HI}}=(\mathcal{T}_{\uparrow}+1)/2,$ we arrive to Eq. (45). We have maximum $\mathcal{F}_{\text{max}}^{\rm HI}=3-2\sqrt{2}\approx 0.17$ at $\mathcal{T}_{\uparrow,\text{ max}}=\sqrt{2}-1$ and, respectively, $R_{\theta}^{\text{max}}=2-\sqrt{2}\approx 0.59$. The total transmission coefficient at this point is given by $\mathcal{T}^{\text{HI}}_{\rm max}=1/\sqrt{2}.$ We notice that Eq. (45) describes both green line in Fig. 10 and dashed black line in Fig. 10.
For almost metallic contact, $t\ll 1$ ($\lambda\to\infty$), and for an arbitrary scattering strength on a magnetic defect, $\mathcal{F}^{\rm HI}$ shows weak oscillations with $\phi$ :
@@
This formula shows weak but sharp resonances with magnetic flux. These resonances are well pronounced in the normalized Fano factor. Indeed, from Eq. (48) for $\theta\ll 1$ we obtain
$\mathcal{F}_{\rm n}^{\text{HI}}\approx\frac{\theta^{2}}{\theta^{2}+\sin^{2}(2\pi\phi)},$
@@ IV Discussion and Conclusion
In the previous sections we analyzed two types of quantum interferometers based on conventional materials and helical edge states. We have demonstrated that Fano factor in both systems is substantially modified by quantum interference effects. It is essential that the interference effects survive in the high temperature limit, $T\gg\Delta.$ The appearance of temperature-robust interference effects stems from trajectories interfering at any energy. In case of symmetric CI for any trajectory entering the interferometer and moving clockwise there is another trajectory moving counterclockwise, and these trajectories interfere destructively at $\phi=1/2.$ Such destructive interference is exact, i.e. the transmission amplitude is identically zero at all energies, $t(\epsilon,\phi=1/2)\equiv 0$ Büttiker et al. (1984) (see also the discussion of the consequences of this identity in Dmitriev et al. (2010, 2015)). As a result, conventional interferometers demonstrate anti-resonances in conductance, $\mathcal{T}^{\rm CI}\to 0$, and resonances in Fano factor, $\mathcal{F}^{\text{CI}}\to 1$, at half-integer flux values [see Eqs. (14) and (15)]. For asymmetric CI there also appear peaks of different height and width at integer flux values. Both $\mathcal{T}^{\rm CI}$ and $\mathcal{F}^{\rm CI}$ are periodic in flux with a period $\Delta\phi=1.$
The peculiarity of interference effects in helical interferometer lies in the fact that they appear only in presence of backscattering defects. In this case trajectories interfering at any energy appear as well, namely, trajectories returning to the defect after $n$ revolutions clockwise and counterclockwise are always interfering. The relative phase of such trajectories is $4\pi n\phi$, and as a result the Aharonov-Bohm oscillation period in HI equals to $\Delta\phi=1/2.$ Thus we have a similarity between the ballistic CI where the scattering at the contacts is always present and HI with backward scatterers.
Interestingly, at $\phi=0$, the dependence of the $\mathcal{F}^{\text{CI}}$ on the transmission coefficient is given by the same formula for the ballistic helical and the conventional interferometers [see Eqs. (22) and (42)]. However, this equation fails for conventional interferometer for $\phi\neq 0$ so that two branches in the dependence $\mathcal{F}^{\text{CI}}(\mathcal{T}^{\text{CI}})$ arise due to the fact that the same conductance can be realized in a conventional interferometer by contacts of different types [see panels (a) and (b) of the Fig. 3].
To conclude, we have developed a theory of shot noise in quantum Aharonov-Bohm interferometers under condition $T\gg\Delta$, Eq. (1), and described in detail two cases: (i) conventional single-channel spinless interferometer and (ii) interferometer based on helical edge states. We demonstrated that in both cases the interference effects strongly modify the shot noise. We have derived general analytical expressions for the Fano factor and demonstrated a resonant behavior of $\mathcal{F}$ on $\phi$ both for CI and HI in the regime of weak tunneling coupling to the leads. The resonant peaks are located at integer and half-integer values of the magnetic flux. An exception is the case of CI with the interferometer arms of equal length, when $\mathcal{F}$ exhibits only resonances at half-integer values of the flux, $\phi=1/2+n$ with integer $n,$ similarly to antiresonances in conductance. The resonant peaks are found to be much more pronounced for CI. In HI even in the regime of well pronounced resonances the relative change in $\mathcal{F}$ in the whole range of $\phi$ is quite small ($\lesssim 0.1$), with resonances having small amplitude and appearing on the top of high background value of $\mathcal{F}$. In both CI and HI, the amplitude of peaks decreases when decreasing the tunneling coupling. For completely closed CI and HI rings, peaks disappear while $\mathcal{F}\to 1/2$. For metallic contacts, $\mathcal{F}$ does not depend on $\phi$ for HI and shows weakly non-linear oscillations with $\phi$ for CI. The magnetic defect strongly enhances the $\mathcal{F}^{\rm HI},$ and there is an optimal value of the defect strength $\theta_{\text{max}}\approx 0.28\,\pi,$ providing the maximum value of the Fano factor, $\mathcal{F}_{\text{max}}^{\rm HI}\approx 0.17.$ On the contrary, for the weak tunneling contacts, the magnetic defect reduces $\mathcal{F}^{\rm HI}$. In the ballistic CI, $\mathcal{F}^{\rm CI}$ also depends on backscattering, defined by the contacts. Similar to HI, there is an optimal value of backscattering, however, in contrast to HI, $\mathcal{F}^{\rm CI}$ has a minimum at this point.
Before closing the paper we make several comments regarding possible experimental measurements of the Fano factor. As for conventional interferometer, it shows different noise for the same transmission coefficient (see, for example, Fig. 7). Hence, measuring, complementary to conductance, also the Fano factor would allow one to determine the type of the tunneling coupling (see Fig. 3a and Fig. 3b) and its strength. In the helical interferometer, measurement of both conductance and the Fano factor would allow one to find probability of backscattering that violate topological protection. Indeed, as seen from Fig. 10 values of $\mathcal{T}^{\rm HI}$ and $\mathcal{F}^{\rm HI}$ unambiguously determines the value of the backscattering probability $R_{\theta}.$ Also, measurement of the flux dependence of the normalized Fano factor and transmission coefficient would allow one to determine type of the contacts. For tunneling contacts and sufficiently weak backscattering we predict sharp resonant behavior of both quantities on the flux (see Fig. 9c), while for the case of “metallic” contact and strong backscattering both Fano factor and conductance would show harmonic Aharonov-Bohm oscillations (see Fig. 9d).
@@ Acknowledgments
The work was carried out with financial support from the Russian Science Foundation (grant No. 20-12-00147-$\Pi$). The work of R.A. Niyazov was also partially supported by the Theoretical Physics and Mathematics Advancement Foundation “BASIS”.
@@ Appendix A Energy averaging
[…]
@@
$\left\langle\frac{1}{(1-\tau_{1}e^{\alpha ikL})}\frac{1}{(1-\tau_{2}e^{\alpha ikL})}\frac{1}{(1-\tau_{3}e^{-\alpha ikL})}\frac{1}{(1-\tau_{4}e^{-\alpha ikL})}\right\rangle_{\varepsilon}$ $=\frac{1-\tau_{1}\tau_{2}\tau_{3}\tau_{4}}{(1-\tau_{1}\tau_{3})(1-\tau_{2}\tau_{3})(1-\tau_{1}\tau_{4})(1-\tau_{2}\tau_{4})}\,,$
@@ Appendix B Conventional interferometer with non-equal arms and non-identical contacts
In this Appendix, we present generalization of the tunneling amplitude $t^{\rm CI}(\varepsilon)$ [see Eq. (17)] for the case of non-identical contacts having different tunneling couplings. The symmetric case is obtained by putting $t_{\text{in,L}}=t_{\text{in,R}}=t_{\text{in}}$, etc. in equations presented below.
$t^{\rm CI}(\varepsilon)=$ $\sum_{n=0}^{\infty}\vec{\alpha}\hat{A}^{n}\vec{\beta}_{0}\,,\quad\vec{\alpha}=\left(\begin{matrix}1\\ 1\end{matrix}\right)\,,\quad\vec{\beta}=\begin{pmatrix}\beta_{0}^{+}\\ \beta_{0}^{-}\end{pmatrix}=t_{\text{in,L}}t_{\text{out,R}}\begin{pmatrix}e^{i(k-2\pi\phi/L)(L/2+a)}\\ e^{i(k+2\pi\phi/L)(L/2-a)}\end{pmatrix}\,,$ (50)
$\hat{A}=$ $e^{ikL}\begin{pmatrix}t_{\text{L}}t_{\text{R}}e^{-i2\pi\phi}+t_{\text{b,L}}t_{\text{b,R}}e^{i2ka}&&t_{\text{L}}t_{\text{b,R}}e^{-i2\pi\phi}+t_{\text{b,L}}t_{\text{R}}e^{i2ka}\\ t_{\text{L}}t_{\text{b,R}}e^{i2\pi\phi}+t_{\text{b,L}}t_{\text{R}}e^{-i2ka}&&t_{\text{L}}t_{\text{R}}e^{i2\pi\phi}+t_{\text{b,L}}t_{\text{b,R}}e^{-i2ka}\end{pmatrix}\,.$
@@
$t(\varepsilon)^{\rm CI}=$ $t_{\text{in,L}}t_{\text{out,R}}e^{-i2\pi\phi a/L}\left(\frac{Ze^{i(kL/2-\pi\phi)}}{1-\tilde{t}^{2}e^{i{kL-2\pi\phi}}}+\frac{Z^{\ast}|_{\text{L}\leftrightarrow\text{R}}e^{i(kL/2+\pi\phi)}}{1-\tilde{t}^{\ast 2}e^{i{kL+2\pi\phi}}}\right)\,,$ (51)
@@
$Z_{1}=$ $Z_{4}-(t_{\text{b,L}}-t_{\text{L}})(t_{\text{b,R}}\sin(2ak)+t_{\text{R}}\sin(2\pi\phi))-i(t_{\text{b,R}}t_{\text{L}}\cos(2ak)+t_{\text{b,L}}t_{\text{R}}\cos(2\pi\phi))\,,$
$Z_{2}=$ $Z_{4}+(t_{\text{b,R}}-t_{\text{R}})(t_{\text{b,L}}\sin(2ak)+t_{\text{L}}\sin(2\pi\phi))-i(t_{\text{b,L}}t_{\text{R}}\cos(2ak)+t_{\text{b,R}}t_{\text{L}}\cos(2\pi\phi))\,,$
$Z_{3}=$ $2Z_{4}\left(t_{\text{b,L}}t_{\text{R}}e^{-i(ak+\pi\phi)}+t_{\text{b,R}}t_{\text{L}}e^{i(ak+\pi\phi)}\right)\,,$
$Z_{4}^{2}=$ $\left(t_{\text{b,L}}^{2}-t_{\text{L}}^{2}\right)\left(t_{\text{b,R}}^{2}-t_{\text{R}}^{2}\right)-(t_{\text{b,L}}t_{\text{b,R}}\cos(2ak)+t_{\text{L}}t_{\text{R}}\cos(2\pi\phi))^{2}\,,$
$\tilde{t}^{2}=$ $iZ_{4}+t_{\text{b,L}}t_{\text{b,R}}\cos(2ak)+t_{\text{L}}t_{\text{R}}\cos(2\pi\phi)\,.$
Assuming that $eV\ll v_{F}/a$, one can replace $ka\rightarrow k_{F}a$ in this equation. Now we see that the expressions $|t^{\rm CI}(\varepsilon)|^{2}$, $|t^{\rm CI}(\varepsilon)|^{4}$ contain the complex quantity $e^{ikL}$ which is regarded as a new variable. Then, one can perform averaging of these expressions over energy by using formulad derived in Appendix  A.

### v2

@@ III.3.3 Limiting cases
$\mathcal{F}^{\mathrm{HI}}=\frac{1}{2}\left[{1-({\mathcal{T}}^{\mathrm{HI}})^{2}}\right].$ (45)
@@
The formulas for the conductance and Fano factor in case of different tunneling contacts can also be obtained in closed form and are presented in C.2. An analog of Eq. (45) in case of different tunneling contacts, in contrast to CI case, cf. Eq. (32) looks a bit more cumbersome
$\mathcal{F}^{\rm HI}=\left[1-({\mathcal{T}}^{\rm HI})^{2}\right]\sqrt{\sinh^{2}(2\beta)+({\mathcal{T}}^{\rm HI})^{2}}$ (46)
$/\left(\sqrt{\sinh^{2}(2\beta)+({\mathcal{T}}^{\rm HI})^{2}}+{\mathcal{T}}^{\rm HI}\cosh(2\beta)\right)\,$,
with the asymmetry parameter $\beta=\frac{1}{2}\ln(t_{\mathrm{R}}/t_{\mathrm{L}})$, Eq. (39).
Figure 10: The combined dependence of Fano factor, $\mathcal{F}^{\mathrm{HI}}$, and the transmission coefficient, $\mathcal{T}^{\mathrm{HI}}$, on the transparency of the contact, $\lambda$, for different strength of the magnetic impurity, $R_{\theta}$, and the magnetic flux $\phi=0.4$, according to Eqs. (40), (41). The black dashed line borders the area of maximum transmission coefficient for given $R_{\theta}$.
@@
$\mathcal{T}^{\rm HI}=\frac{1}{2}{\tanh[2\lambda]},\quad\mathcal{F}^{\rm HI}=\frac{1}{2\cosh^{2}[2\lambda]},$ (47)
$\mathcal{F}^{\rm HI}=\frac{1}{2}-2\left(\mathcal{T}^{\rm HI}\right)^{2}.$ (48)
@@
$\mathcal{T}^{\mathrm{HI}}=1-\frac{R_{\theta}}{2},$ (49)
$\mathcal{F}^{\mathrm{HI}}=R_{\theta}\frac{1-R_{\theta}}{2-R_{\theta}}=\frac{(2\mathcal{T}^{\mathrm{HI}}-1)(1-\mathcal{T}^{\mathrm{HI}})}{\mathcal{T}^{\mathrm{HI}}}\,.$
@@
Equation (49) has a simple physical meaning. For $r=1,$ the windings are absent, and there are two parallel conductors, i.e. lower and upper arms of the HI. The lower arm, with spin $\downarrow$, is an ideal conductor with the transmission coefficient $\mathcal{T}_{\downarrow}=1$ while the transmission through the upper arm is controlled by the magnetic defect, $\mathcal{T}_{\uparrow}=\cos^{2}\theta,$ varying from 0 to $1.$ Then, we get
@@
Taking into account that the total transmission coefficient in this case is given by $\mathcal{T}^{\mathrm{HI}}=(\mathcal{T}_{\uparrow}+1)/2,$ we arrive to Eq. (49). We have maximum $\mathcal{F}_{\mathrm{max}}^{\rm HI}=3-2\sqrt{2}\approx 0.17$ at $\mathcal{T}_{\uparrow,\mathrm{max}}=\sqrt{2}-1$ and, respectively, $R_{\theta}^{\mathrm{max}}=2-\sqrt{2}\approx 0.59$. The total transmission coefficient at this point is given by $\mathcal{T}^{\mathrm{HI}}_{\rm max}=1/\sqrt{2}.$ We notice that Eq. (49) describes both black line in Fig. 8a and dashed black line in Fig. 10.
In order, to combine here analytical formulas for all relevant cases, we also present at the end of this section several equations obtained previously in Ref. Niyazov et al. (2024). For almost metallic contact, $t\ll 1$ ($\lambda\to\infty$), and for an arbitrary scattering strength on a magnetic defect, $\mathcal{F}^{\rm HI}$ shows weak oscillations with $\phi$:
@@
This formula shows weak but sharp resonances with magnetic flux. These resonances are well pronounced in the normalized Fano factor. Indeed, from Eqs.  (43) and (52) for $\theta\ll 1$ we obtain
$\mathcal{F}_{\rm n}^{\mathrm{HI}}\approx\frac{\theta^{2}}{\theta^{2}+\sin^{2}(2\pi\phi)},$
@@ IV Conclusion
In the previous sections we analyzed two types of quantum interferometers based on conventional materials and helical edge states. We have demonstrated that the shot noise in both systems is substantially modified by quantum interference effects. It is essential that the interference effects survive in the high temperature limit, $T\gg\Delta.$ The appearance of temperature-robust interference effects stems from trajectories interfering at any energy. In case of symmetric CI for any trajectory entering the interferometer and moving clockwise there is another trajectory moving counterclockwise, and these trajectories interfere destructively at $\phi=1/2.$ Such destructive interference is exact, i.e. the transmission amplitude is identically zero at all energies, $t(\epsilon,\phi=1/2)\equiv 0$ Büttiker et al. (1984) (see also the discussion of the consequences of this identity in Dmitriev et al. (2010, 2015)). The peculiarity of interference effects in HI lies in the fact that they appear only in presence of backscattering defects. In this case trajectories interfering at any energy appear as well, namely, trajectories returning to the defect after $n$ revolutions clockwise and counterclockwise are always interfering. The relative phase of such trajectories is $4\pi n\phi$, and as a result the Aharonov-Bohm oscillation period in HI equals to $\Delta\phi=1/2,$ in contrast to CI, where the oscillation period is equal to $\Delta\phi=1.$
To conclude, in this paper we discuss and compare noise characteristics of two types of Aharonov-Bohm interferometers, based on conventional materials and on helical edges states of TI. We focus on the case of relatively high temperatures, which partially smear the interference effects, all the while leaving many remarkable quantum features intact. In particular, we show that even in case of higher temperatures the noise Fano factor reveals a characteristic periodicity in the magnetic field. We demonstrate different patterns of this periodicity with resonances and anti-resonances at integer and half-integer values of the flux (see Fig. 6 for CI and Fig. 9 for HI).
One of the most important conclusions following from our calculations is that the interference effects are much more pronounced for CI and, consequently, the influence of the magnetic field on the Fano factor is much stronger as compared to HI. The reason is that the interference comes into play for CI already in the ballistic case, when impurities in the arms of CI are absent, while the Aharonov-Bohm peaks in HI are proportional to the backscattering probability.
We show that asymmetrical placements of contacts results in strong modification of noise Fano factor in CI in contrast to HI. In particular, as seen from Eqs. (30) and (31), the asymmetry of the CI strongly affects the noise intensity even for fixed magnetic flux, so that changing difference of shoulder’s lengths, $a,$ within a small interval about the Fermi wavelength, we change $\mathcal{F}_{\rm CI}$ by a factor on the order of unity. We also find that the characteristics of the contact between the CI and the leads strongly modify the noise in this case, even in the ballistic regime inside the CI. In turn, the noise in case of HI is strongly impacted by the backscattering defects inside the HI ring.
We also demonstrate that the simultaneous measurement of conductance and Fano factor in HI allows one to determine the strength of backscattering defects that violate topological protection in the realistic setups. As for CI, it can show different noise for the same transmission coefficient (see, for example, Fig. 7). Hence, measuring, complementary to conductance, also the Fano factor of CI would help to experimentally determine the type of the contacts to leads (see Fig. 3a and Fig. 3b) and the strength of tunneling coupling.
Data availability statement
All data that support the findings of this study are included within the article (and any supplementary files).
@@ Acknowledgments
The work was carried out with financial support from the Russian Science Foundation (grant No. 25-12-00212), https://rscf.ru/en/project/25-12-00212/.
@@ Appendix A Energy averaging
Throughout the paper we average observable quantities over a small temperature window in the vicinity of the Fermi level in the limit $T\gg\Delta$. For the linearized form of the spectrum, with $\varepsilon=v_{F}k$, the energy averaging is reduced at high temperatures to calculating the integral $\left\langle\cdots\right\rangle_{\varepsilon}=\Delta^{-1}\int_{0}^{\Delta}d\varepsilon\,(\cdots)=\frac{L}{2\pi}\int_{0}^{2\pi/L}dk\,(\cdots)$. The averaging over $kL$, is reduced to integration over the unit circle $z=e^{ikL}$ in complex plane and is easily performed by residues. The formulas which were used for calculations are listed below:
@@
$\left\langle\frac{1}{(1-\tau_{1}e^{\alpha ikL})}\frac{1}{(1-\tau_{2}e^{\alpha ikL})}\frac{1}{(1-\tau_{3}e^{-\alpha ikL})}\frac{1}{(1-\tau_{4}e^{-\alpha ikL})}\right\rangle_{\varepsilon}$
$=\frac{1-\tau_{1}\tau_{2}\tau_{3}\tau_{4}}{(1-\tau_{1}\tau_{3})(1-\tau_{2}\tau_{3})(1-\tau_{1}\tau_{4})(1-\tau_{2}\tau_{4})}\,$,
@@ Appendix B Conventional interferometer with non-equal arms and non-identical contacts
In this Appendix, we present generalization of the tunneling amplitude $t^{\rm CI}(\varepsilon)$ [see Eq. (19)] for the case of non-identical contacts having different tunneling couplings. The symmetric case is obtained by putting $t_{\mathrm{in,L}}=t_{\mathrm{in,R}}=t_{\mathrm{in}}$, etc. in equations presented below.
$t^{\mathrm{CI}}(\varepsilon)=$ $\sum_{n=0}^{\infty}\vec{\alpha}\hat{A}^{n}\vec{\beta}_{0}\,,\quad\vec{\alpha}=\left(\begin{array}[]{c}1\\ 1\end{array}\right)\,,\quad\vec{\beta}=\left(\begin{array}[]{c}\beta_{0}^{+}\\ \beta_{0}^{-}\end{array}\right)=t_{\mathrm{in,L}}t_{\mathrm{out,R}}\left(\begin{array}[]{c}e^{i(k-2\pi\phi/L)(L/2+a)}\\ e^{i(k+2\pi\phi/L)(L/2-a)}\end{array}\right)\,$, (54)
$\hat{A}$= $e^{ikL}\left(\begin{array}[]{cc}t_{\mathrm{L}}t_{\mathrm{R}}e^{-i2\pi\phi}+t_{\mathrm{b,L}}t_{\mathrm{b,R}}e^{i2ka}&t_{\mathrm{L}}t_{\mathrm{b,R}}e^{-i2\pi\phi}+t_{\mathrm{b,L}}t_{\mathrm{R}}e^{i2ka}\\ t_{\mathrm{L}}t_{\mathrm{b,R}}e^{i2\pi\phi}+t_{\mathrm{b,L}}t_{\mathrm{R}}e^{-i2ka}&t_{\mathrm{L}}t_{\mathrm{R}}e^{i2\pi\phi}+t_{\mathrm{b,L}}t_{\mathrm{b,R}}e^{-i2ka}\end{array}\right)\,.$
@@
$t(\varepsilon)^{\rm CI}=$ $t_{\mathrm{in,L}}t_{\mathrm{out,R}}e^{-i2\pi\phi a/L}\left(\frac{Ze^{i(kL/2-\pi\phi)}}{1-\tilde{t}^{2}e^{i{kL-2\pi\phi}}}+\frac{Z^{\ast}|_{\mathrm{L}\leftrightarrow\mathrm{R}}e^{i(kL/2+\pi\phi)}}{1-\tilde{t}^{\ast 2}e^{i{kL+2\pi\phi}}}\right)\,$, (55)
@@
$Z_{1}=$ $Z_{4}-(t_{\mathrm{b,L}}-t_{\mathrm{L}})(t_{\mathrm{b,R}}\sin(2ak)+t_{\mathrm{R}}\sin(2\pi\phi))-i(t_{\mathrm{b,R}}t_{\mathrm{L}}\cos(2ak)+t_{\mathrm{b,L}}t_{\mathrm{R}}\cos(2\pi\phi))\,$,
$Z_{2}=$ $Z_{4}+(t_{\mathrm{b,R}}-t_{\mathrm{R}})(t_{\mathrm{b,L}}\sin(2ak)+t_{\mathrm{L}}\sin(2\pi\phi))-i(t_{\mathrm{b,L}}t_{\mathrm{R}}\cos(2ak)+t_{\mathrm{b,R}}t_{\mathrm{L}}\cos(2\pi\phi))\,$,
$Z_{3}=$ $2Z_{4}\left(t_{\mathrm{b,L}}t_{\mathrm{R}}e^{-i(ak+\pi\phi)}+t_{\mathrm{b,R}}t_{\mathrm{L}}e^{i(ak+\pi\phi)}\right)\,$,
$Z_{4}^{2}=$ $\left(t_{\mathrm{b,L}}^{2}-t_{\mathrm{L}}^{2}\right)\left(t_{\mathrm{b,R}}^{2}-t_{\mathrm{R}}^{2}\right)-(t_{\mathrm{b,L}}t_{\mathrm{b,R}}\cos(2ak)+t_{\mathrm{L}}t_{\mathrm{R}}\cos(2\pi\phi))^{2}\,$,
$\tilde{t}^{2}=$ $iZ_{4}+t_{\mathrm{b,L}}t_{\mathrm{b,R}}\cos(2ak)+t_{\mathrm{L}}t_{\mathrm{R}}\cos(2\pi\phi)\,.$
Assuming that $eV\ll v_{F}/a$, one can replace $ka\rightarrow k_{F}a$ in this equation. Now we see that the expressions $|t^{\rm CI}(\varepsilon)|^{2}$, $|t^{\rm CI}(\varepsilon)|^{4}$ contain the complex quantity $e^{ikL}$ which is regarded as a new variable. Then, one can perform averaging of these expressions over energy by using formulas derived in A.